OpenStudy (anonymous):

What is the area of an equilateral triangle with side lengths of 8 cm?

OpenStudy (anonymous):

We need to use Pythagoras here. to find the height. The height will be right down the middle. Half of 8 = 4. If (a^2)+(b^2)=c^2 You know a^2 is 4 and c^2 is 8. b^2= c^2-a^2 b=8−4−−−−√ 8−4−−−−√=4√=2 2 is the height. Using the formula a=bh/2 b= 8 h= 2 8*2=16 16/2=8 therefore 8 is the area of this equilateral triangle.

OpenStudy (anonymous):

thank you

terenzreignz (terenzreignz):

No.

terenzreignz (terenzreignz):

@KaylaRose98 the formula for the area of an equilateral triangle with side equal to a is \[\huge \frac{a^2\sqrt3}{4}\]

OpenStudy (anonymous):

27.713 then?

terenzreignz (terenzreignz):

Whatever that might be. I'd leave it with the \(\sqrt 3\) but if the teacher wants a decimal answer, then go for it

OpenStudy (anonymous):

thank yo

OpenStudy (anonymous):

you

OpenStudy (anonymous):

can you answer my question please?

terenzreignz (terenzreignz):

Might as well derive it, while you guys are watching... 

terenzreignz (terenzreignz):



terenzreignz (terenzreignz):

Using the Pythagorean theorem, the height is given by \[\huge \sqrt{a^2 - \frac{a^2}{4}} =\sqrt{a^2\left(1-\frac14\right)}=a\sqrt{\frac34}=\frac{a\sqrt3}{2}\]

terenzreignz (terenzreignz):

Now, since \[h = \frac{a\sqrt3}{2}\]and the base\[b = a\] Then the area \[\huge \frac{bh}2=\frac{a\times \frac{a\sqrt3}{2}}{2}=\frac{a^2\sqrt3}{4}\]

OpenStudy (dwelsher):

~27.72sq cm